a solution has a POH at 6.50 at 50° C what is the Ph of the solution given that K=6.22 X 10-15

Use pOH in calculations Question A solution has a pOH of 6.50 at 50°C. What is the pH of the solution given that Kw 6.22 x 10 please explain HOW too
Answer #1

We know that

pH + pOH = pKw

Given that

pOH = 6.50

Kw = 6.22 × 10–15

pKw = – logKw = – log(6.22 × 10–15) = 14.20621 = 14.21

pH = pKw – pOH = 14.21 – 6.50 = 7.71

pH of the solution = 7.71